OpenStudy (anonymous):

Find the percent of markdown on a skirt that is regularly priced at $50 and is on sale for $40.

OpenStudy (anonymous):

You had something that costed 50 and now is costing 40. So all you want to know is the amount X that is multiplied by 50 to give 40. Put it in an equation: 50* x = 40 can you continue from here?

OpenStudy (anonymous):

well 40 = 50(100% - markdown %)

OpenStudy (anonymous):

since 40 = 50 - 50*(markdown %)

OpenStudy (anonymous):

Yeah @ Diogo

OpenStudy (anonymous):

Thanks @ RONNCC